Solve for​ x: 0.25x + .01(x−3)=1.1

Show work
Answer #1

SOLUTION;;

The equation is

0.25x+0.01(x-3)=1.1

0.25x+0.01x-0.03=1.1

0.25x+0.01x=0.03+1.1

0.26x=1.13

x=\frac{1.13}{0.26}

x=\frac{113}{26}

Answer

x=\frac{113}{26}